OpenStudy (anonymous):

assing through (-3, -7) and parallel to the line whose equation is y=-4x+5

OpenStudy (anonymous):

Lines are parallel means their slopes are equal As line y=-4x+5 has slope -4, that will be the slope of second line as well Now if a line is passing a given poin (x1,y1) & having a slope m then its equation is given by y-y1= m(x-x1) Here (x1,y1) is (-3,-7) & m=-4 Now u can proceed from here

OpenStudy (dan815):

y=-4x+c where c can be any number are all the lines parallel to that given equation, u need to know which one of these linnes passes thru that given point